# A Syntax-Injected Approach for Faster and More Accurate Sentiment Analysis

### v1

@@ Abstract
Sentiment Analysis (SA) is a crucial aspect of Natural Language Processing (NLP), addressing subjective assessments in textual content. Syntactic parsing is useful in SA because explicit syntactic information can improve accuracy while providing explainability, but it tends to be a computational bottleneck in practice due to the slowness of parsing algorithms. This paper addresses said bottleneck by using a SEquence Labeling Syntactic Parser (SELSP) to inject syntax into SA. By treating dependency parsing as a sequence labeling problem, we greatly enhance the speed of syntax-based SA. SELSP is trained and evaluated on a ternary polarity classification task, demonstrating its faster performance and better accuracy in polarity prediction tasks compared to conventional parsers like Stanza and to heuristic approaches that use shallow syntactic rules for SA like VADER. This increased speed and improved accuracy make SELSP particularly appealing to SA practitioners in both research and industry. In addition, we test several sentiment dictionaries on our SELSP to see which one improves the performance in polarity prediction tasks. Moreover, we compare the SELSP with Transformer-based models trained on a 5-label classification task. The results show that dictionaries that capture polarity judgment variation provide better results than dictionaries that ignore polarity judgment variation. Moreover, we show that SELSP is considerably faster than Transformer-based models in polarity prediction tasks.
A Syntax-Injected Approach for Faster and More Accurate Sentiment Analysis
Muhammad Imran thanks:   Muhammad Imran and Olga Kellert contributed equally to this work.    Corresponding author: Muhammad Imran (m.imran@udc.es), Olga Kellert  and Carlos Gómez-Rodríguez Grupo LyS, CITIC, Depto. de Ciencias de la Computación y Tecnologías de la Información, Universidade da Coruña, Campus de Elviña s/n, 15071 A Coruña, Spain
1 Introduction
@@
Sentiment Analysis (SA) holds a pivotal role in Natural Language Processing (NLP), dealing with the identification of subjective assessments, such as opinions on hotels in textual content (subjectivity identification task), and predicting the polarity of these assessments, whether they are positive or negative (Polarity Classification task, PC), as exemplified in (1) (Cui et al., 2023).
(1) Anne: This hotel is awesome. (PC: Positive)
Ben: This hotel is terrible. (PC: Negative)
@@
Some of the approaches to SA are based on locating sentiment-bearing words on sentences, like “awesome” and “terrible” in (1), and using them as main polarity predictors. These are called sentiment words, and are listed together with their associated polarities in sentiment dictionaries like SO-CAL (Brooke et al., 2009). These approaches to SA provide an intuitive, transparent, and easily explainable analysis; as the sentiment polarity associated with a sentence can be traced back to the specific words that were used to deduce it. This is in contrast to supervised approaches that learn predictions automatically, using black-box models that do not provide any such transparency (Cui et al., 2023).
@@
However, the polarity of a sentiment word can be shifted by various linguistic elements, notably including negation, as in "This hotel is not good". As a consequence, if high accuracy is desired, dictionary-based approaches cannot simply rely on sentiment dictionaries like SO-CAL to perform SA tasks. These approaches should also somehow model the relation between sentiment words and polarity shifting elements like negation. Since the scope of relevant phenomena such as negation can be determined from sentence syntax, syntactic parsing is a natural way to deal with the syntactic relations between sentiment words and polarity-shifting constructs. This is done via rules that predict sentence polarity from a parse tree in an explainable and transparent way, while being more accurate than simple local heuristics like e.g. having the negation affect a fixed-length window of tokens (Thelwall et al., 2010). The most common formalism used for syntactic parsing is Universal Dependencies (UD) (Zeman et al., 2023). This is a dependency parsing formalism, i.e., it represents the syntax of a sentence with a parse tree composed of a set of directed binary relations between words called dependencies, which describe the syntactic roles (e.g., direct object, subject, adverbial modifier) that participate in the sentence. Stanza is an example of a popular NLP tool that provides syntactic dependency parsing in the UD formalism in several languages, including Spanish and English (Qi et al., 2020). It has been shown that approaches that use syntactic parsing for an explainable and transparent SA based on the UD formalism or predecessor versions (Vilares et al., 2017; Kellert et al., 2023) show better performance in polarity prediction tasks than heuristic approaches that do not rely on syntactic parsing at all such as VADER (Hutto and Gilbert, 2014), as demonstrated by Gómez-Rodríguez et al. (2017).
@@
Despite the clear advantages of using syntactic parsing for SA, conventional syntactic parsers like Stanza (Qi et al., 2020) are often slow and computationally costly (Gómez-Rodríguez and Vilares, 2018; Roca et al., 2023; Strzyz et al., 2019b). Several solutions have been suggested in recent years to address this problem by developing faster systems by the transformation of dependency parsing into a SEquence Labeling (SEL) task (Gómez-Rodríguez and Vilares, 2018; Roca et al., 2023; Strzyz et al., 2019b). However, so far, SEL parsers have not been applied to SA. This paper fills this gap. Thus, we adopt the SEL dependency parsing approach for sentiment analysis, to speed up the polarity prediction task and evaluate its practicality. In addition, we test our approach on several dictionaries to see which one improves the accuracy of the SEL-based parser.
@@
Our contributions are: (1) We present, to the best of our knowledge, the first approach that performs fast, syntax-based polarity classification using sequence labeling parsing. For this purpose, we first train an SEL system to perform syntactic dependency parsing (SELSP) in English and Spanish. After measuring its parsing accuracy on test sets of these languages, we then apply our SESLP in the domain of SA for the polarity prediction task, using syntax-based rules to calculate polarity in English and Spanish. (2) We compare the performance of our approach in terms of runtime and accuracy in both languages to using the same syntactic rules on top of the well-known dependency parser from Stanza, as well as to VADER, a heuristic approach to SA that does not use syntax trees and is widely used in industry, and to a state-of-the-art Transformer-based (RoBERTa) model. (3) We study the influence of sentiment dictionaries and their interaction with syntax-based systems by using several sentiment dictionaries, and combinations thereof, on top of the same set of rules and the two mentioned parsers.
2 Related Work
@@
2.1 ML/DL methods and hybrid methods
Supervised approaches using machine learning, including deep learning architectures like LSTM or CNN and pre-trained models like BERT and RoBERTa, are a common practice in SA (see Cui et al. (2023)). These approaches use the standard supervised learning methodology where a training dataset is used during the training process to learn correlations between the specific input text and the sentiment polarity, and the resulting trained system is then evaluated on a separate test dataset and/or used in production. The performance of ML/DL-based approaches is known to be relatively high if the learning and prediction tasks are performed on a similar dataset or corpus, as they adapt especially well to the particularities of the training dataset. This means that they often dominate the rankings in shared tasks, such as that of Rest-Mex 2023 Álvarez Carmona et al. (2023), where training and test datasets are obtained from the same source and exhibit similar characteristics in terms of topic, length, style, etc. However, their performance can degrade on texts that differ from the training dataset, even within the same domain (see e.g. Table 5 of Vilares et al. (2017)).
@@ 2.1 ML/DL methods and hybrid methods
Some approaches, like (Yang et al., 2020), use sentiment dictionaries to provide or modify the input information fed to supervised machine learning approaches. Thus, they are often considered hybrid approaches Cui et al. (2023) between ML/DL methods and the lexicon-based methods described in the next section. While exploiting sentiment dictionaries is useful to improve accuracy, their machine-learning-based core means that they still share the drawbacks described above.
2.2 Lexicon-based methods
@@
The lexicon-based methodology makes use of a sentiment dictionary that contains sentiment words with corresponding polarity scores to find out the overall opinion or polarity of a sentence or text Taboada et al. (2011); Hutto and Gilbert (2014); Vilares et al. (2017); Kellert et al. (2023). These approaches fall into various branches depending on how they deal with syntactic rules that can shift or intensify the polarity. Non-compositional and non-recursive rule-based approaches such as SO-CAL (Taboada et al., 2011), SentiStrength (Thelwall et al., 2012) or VADER Hutto and Gilbert (2014) use heuristic rules to account for polarity changes due to negation, modification, or other syntactic processes. Other unsupervised approaches use a more linguistically-informed architecture to account for SA by exploiting general syntactic rules that influence sentiment polarity, requiring a parse tree Tan et al. (2012); Vilares et al. (2017); Kellert et al. (2023).
@@
Lexicon-based methods are considered to be simple to use and cost-effective as there is no need to train on huge amounts of data, and they work well across domains. For these reasons, in spite of less attention in the NLP literature in recent years due to the general trend towards machine learning in the field, they still tend to be the most commonly used in industry and production domains such as biomedicine (see e.g. Table 5 in (Denecke and Reichenpfader, 2023)) or economics (Algaba et al., 2020).
@@
To maximize accuracy, lexicon-based methods need to be able to reliably determine the scope of various relevant syntactic phenomena such as negation or intensification. As mentioned above, the most basic form of scope detection is to use simple local heuristics such as a fixed-length window, but this has limitations in terms of accuracy. For example, systems like SentiStrength or VADER can detect that “good” is negated in “not good”, but not in “not in any respect good”, due to the long-distance dependency. Instead, one can use machine learning classifiers to determine scope (see (Cruz et al., 2016; Jiménez-Zafra et al., 2021) and references therein), but this entails the disadvantages previously discussed for ML approaches, including the need for specific data and lack of explainability. Thus, the alternative on which we focus our proposal is to use syntactic rules that operate on a parse tree (Jia et al., 2009; Vilares et al., 2015; Jimenez Zafra et al., 2019; Vilares et al., 2017; Kellert et al., 2023). The next sections address methods for syntactic parsing and how they can be used for polarity classification.
2.3 Syntactic parsing for sentiment analysis
@@
Syntactic parsing is the task of determining the hierarchical structure of a sentence in natural language. For this purpose, two main kinds of representations are used in NLP: constituent parsing, where the sentence is recursively divided into smaller units, and dependency parsing, which focuses on representing relations between words. While constituent parsing has seen some use in sentiment analysis (Socher et al., 2013), dependency parsing is the dominant alternative for the task, as it straightforwardly represents what words are modified by other words. In particular, it is the basis of all the methods based on syntactic rules cited in the previous section.
@@
In the last few years, Universal Dependencies (UD) (Zeman et al., 2023) has become the predominant framework for dependency parsing, especially in multilingual setups, as it provides harmonized annotation guidelines and freely-available corpora for over 100 diverse languages.
@@
Kellert et al. (2023) make use of UD-based syntactic parsing and exploit dependency relations expressed in UD between sentiment words and other words that can shift, strengthen, or weaken the polarity of sentiment words (see also Vilares et al. (2017) for application of syntactic parsing for SA using a predecessor framework to UD). The first stage of the approach in (Kellert et al., 2023) is to search for sentiment words taken from a sentiment dictionary SO-CAL (Brooke et al., 2009) in the input text, such as the sentiment word wonderful (in Figure 2). Then, the system traverses the dependency tree and applies a set of syntax-based rules to check whether negation words or other modifiers and relevant syntactic phenomena change the polarity of sentiment words (see the dependency tree in Figure 2). This syntactic approach applies recursive and compositional rules to negation and intensification, to identify the right scope of polarity shifting elements and to properly compute the sentence polarity (Vilares et al., 2017; Kellert et al., 2023). This means that the computing of the sentence polarity is done bottom-up: first, intensifiers and polarity-shifting syntactic phenomena in the structurally lowest levels of the tree are used to modify the polarity of their modified sentiment-bearing words and expressions. Then, structurally higher polarity shifting elements are computed, and polarities from converging branches of the dependency tree are aggregated, until the root of the sentence structure is reached and the final combined polarity for the whole sentence is obtained (Vilares et al., 2017; Kellert et al., 2023). For instance, in the sentence in Figure 2, first, the modification of the adjective “kind” by the adverb “very” is computed, and then the conjunction between the sentiment word “wonderful” and the output of “very kind” is computed. Since there are no further polarity-shifting elements, the polarity of this conjunction is output as sentence polarity.
2.4 Dependency parsing approaches
@@
To successfully use UD-based dependency parsing for sentiment analysis as described in Section 2.3, one ideally would want the parser to be accurate, so the obtained trees provide reliable information to compute polarity, and efficient, so that the system can operate at a large scale with a reasonable resource consumption. However, choosing a parsing approach involves a tradeoff between speed and accuracy, as the most accurate systems tend to be slower (Anderson and Gómez-Rodríguez, 2021). In this respect, it is worth noting that Gómez-Rodríguez et al. (2017) showed that the polarity classification task has modest accuracy requirements, i.e., there is a “good enough” parsing accuracy threshold past which increasing accuracy doesn’t bring further improvements to the end task. On the other hand, speed is often a practical bottleneck, especially if a large number of messages need to be processed.
@@
However, as already noted in § Introduction, conventional syntactic parsers based on ad-hoc algorithms, like the one included in the Stanza library (Qi et al., 2020) are often slow and computationally costly (Gómez-Rodríguez and Vilares, 2018; Roca et al., 2023; Strzyz et al., 2019b). However, in recent years, an alternative approach has been developed that provides considerably faster parsing, based on viewing dependency parsing as a sequence labeling task. Sequence labeling can be performed in linear time with respect to the length of the input, and is highly parallelizable, so it provides high practical speeds (Strzyz et al., 2019b; Anderson and Gómez-Rodríguez, 2021).
@@
Previous work on syntactic parsing as sequence labeling has been applied to several NLP tasks such as integration of constituency and dependency parsing (Strzyz et al., 2019a), semantic role labeling (Wang et al., 2019) or LLM probing (Vilares et al., 2020), where its speed and simplicity can be useful. CoDeLin (COnstituent and DEpendency LINearization system, Roca et al. (2023)) is a recent comprehensive library for parsing as sequence labeling that supports both constituent and dependency parsing. The basic idea behind dependency parsing in CoDeLIn is to assign to each word a tag with information about a part of the dependency tree involving that word, in such a way that a full tree can be decoded from the sequence of labels for a given sentence. For this purpose, there are several types of encoding functions, each of which corresponds to a different way of representing the dependencies between words in a sentence in the form of tags associated to each of the words. We now summarize the framework and some of the encodings of interest. More information, including further encoding functions, can be found in (Strzyz et al., 2019b; Roca et al., 2023; Gómez-Rodríguez et al., 2023).
@@
Following (Roca et al., 2023), let $T_{|w|}$ be the set of dependency trees for sentences of length $|w|$. Each element of $T_{|w|}$ is a dependency tree $t_{w}=(V,A)$ for a sentence $w=[w_{1},...,w_{|w|}]$ where each $i\in V$ denotes the position of the word $w_{i}$, and $A$ is the set of dependency relations between words, where each $a\in A$ is of the form $a=(h,r,d)$ such that $h\in V$ is the index of the head of the relation, $d\in V$ corresponds to the index of the dependent, and $r\in R$ is a label describing a syntactic function, i.e. the type of syntactic relation between head and dependent (such as subject, modifier, indirect object, etc.). We can cast the dependency parsing task as a sequence labeling task by defining a set of labels $L$ that will let us encode each tree $t_{w}\in T_{|w|}$ as a sequence of $|w|$ labels $l_{w}\in L^{|w|}$, and an encoding function $F_{|w|}:T_{|w|}\rightarrow L^{|w|}$ that allows us to translate the trees into those labels. $F_{|w|}$ must be injective, so that each tree is encoded as a unique label sequence and correct label sequences can be decoded back into trees. While injectivity is not enough to guarantee that incorrect label sequences that might be output by a machine learning sequence-labeling system can also be decoded (that would require surjectivity, which known encoding functions do not have), such sequences are handled in practice with straightforward heuristics, like ignoring labels that are contradictory or violate tree constraints.
@@
Figure 1 illustrates how a sentence’s parse tree is represented in terms of different encodings in CoDeLIn (Strzyz et al., 2019b; Roca et al., 2023). Once one has chosen a tree encoding, training a parser is straightforward: a sequence-labeling model is trained to label sentences with the syntactic labels in the selected encoding, and at test time, the generated labels are decoded back to trees by using the displacements (and PoS tags, in the PoS-based encoding) to find the head of each node and build the corresponding dependency. If the model outputs a wrong label that causes a contradiction or constraint violation (e.g. a head offset pointing to a non-existing word position like position 6 in a sentence of length 4, or causing a cycle), that label is simply ignored and no dependency is created.
Figure 1: A short version for demonstration of Sequence Labeling from (Strzyz et al., 2019b).
@@ 3.1 Data
We are using different datasets for the different tasks in our pipeline. For dependency parsing, we use the datasets UD EWT English (Silveira et al., 2014) and UD Spanish AnCora (Taulé et al., 2008) with standard splits.

### v2

@@ Abstract
Sentiment Analysis (SA) is a crucial aspect of Natural Language Processing (NLP), addressing subjective assessments in textual content. Syntactic parsing is useful in SA because explicit syntactic information can improve accuracy while providing explainability, but it tends to be a computational bottleneck in practice due to the slowness of parsing algorithms. This paper addresses said bottleneck by using a SEquence Labeling Syntactic Parser (SELSP) to inject syntax into SA through a sentiment analysis engine that relies on syntactic rules. By treating dependency parsing as a sequence labeling problem, we greatly enhance the speed of syntax-based SA. SELSP is trained and evaluated on a ternary polarity classification task, demonstrating its faster performance and better accuracy in polarity prediction tasks compared to conventional parsers like Stanza and to heuristic approaches that use shallow syntactic rules for SA like VADER. This increased speed and improved accuracy make SELSP particularly appealing to SA practitioners in both research and industry. In addition, we test several sentiment dictionaries on our SELSP to see which one provides the best performance in polarity prediction tasks. Moreover, we compare the SELSP with Transformer-based models trained on a 5-label classification task. The results show that dictionaries that capture polarity judgment variation provide better results than dictionaries that ignore polarity judgment variation. Furthermore, we show that SELSP is considerably faster than Transformer-based models in polarity prediction tasks.
Introduction
@@
Sentiment Analysis (SA) holds a pivotal role in Natural Language Processing (NLP), dealing with the identification of subjective assessments, such as opinions on hotels in textual content (subjectivity identification task), and predicting the polarity of these assessments, whether they are positive or negative (Polarity Classification task, PC), as exemplified in (1) (Cui et al.,, 2023).
(1) Anne: This hotel is awesome. (PC: Positive) Ben: This hotel is terrible. (PC: Negative)
@@
Some of the approaches to SA are based on locating sentiment-bearing words on sentences, like “awesome” and “terrible” in (1), and using them as main polarity predictors. These are called sentiment words, and are listed together with their associated polarities in sentiment dictionaries like SO-CAL (Brooke et al.,, 2009). These approaches to SA provide an intuitive, transparent, and easily explainable analysis; as the sentiment polarity associated with a sentence can be traced back to the specific words that were used to deduce it. This is in contrast to supervised approaches that learn predictions automatically, using black-box models that do not provide any such transparency (Cui et al.,, 2023).
@@
However, the polarity of a sentiment word can be shifted by various linguistic elements, notably including negation, as in “This hotel is not good”. As a consequence, if high accuracy is desired, dictionary-based approaches cannot simply rely on sentiment dictionaries like SO-CAL to perform SA tasks. These approaches should also somehow model the relation between sentiment words and polarity shifting elements like negation. Since the scope of relevant phenomena such as negation can be determined from sentence syntax, syntactic parsing is a natural way to deal with the syntactic relations between sentiment words and polarity-shifting constructs. This is done via rules that predict sentence polarity from a parse tree in an explainable and transparent way, while being more accurate than simple local heuristics like e.g. having the negation affect a fixed-length window of tokens (Thelwall et al.,, 2010). The most common formalism used for syntactic parsing is Universal Dependencies (UD) (Zeman et al.,, 2023).
This is a dependency parsing formalism, i.e., it represents the syntax of a sentence with a parse tree composed of a set of directed binary relations between words called dependencies, which describe the syntactic roles (e.g., direct object, subject, adverbial modifier) that participate in the sentence. Stanza is an example of a popular NLP tool that provides syntactic dependency parsing in the UD formalism in several languages, including Spanish and English, our focus languages in this paper (Qi et al.,, 2020). It has been shown that approaches that use syntactic parsing for an explainable and transparent SA based on the UD formalism or predecessor versions (Vilares et al.,, 2017; Kellert et al.,, 2023) show better performance in polarity prediction tasks than heuristic approaches that do not rely on syntactic parsing at all such as VADER (Hutto and Gilbert,, 2014), as demonstrated by Gómez-Rodríguez et al., (2019).
@@
Despite the clear advantages of using syntactic parsing for SA, conventional syntactic parsers like Stanza (Qi et al.,, 2020) are often slow and computationally costly (Gómez-Rodríguez and Vilares,, 2018; Roca et al.,, 2023; Strzyz et al., 2019b, ). Several solutions have been suggested in recent years to address this problem by developing faster systems by the transformation of dependency parsing into a SEquence Labeling (SEL) task (Gómez-Rodríguez and Vilares,, 2018; Roca et al.,, 2023; Strzyz et al., 2019b, ). However, so far, SEL parsers have not been applied to SA. This paper fills this gap. Thus, we adopt the SEL dependency parsing approach for sentiment analysis, to speed up the application of syntactic rules for polarity prediction and evaluate its practicality. In addition, we test our approach with several dictionaries to see which one yields the best accuracy when used with the SEL-based system.
@@
Our contributions are: (1) We present, to the best of our knowledge, the first approach that performs fast, syntax-based polarity classification using sequence labeling parsing. For this purpose, we first train an SEL system to perform syntactic dependency parsing (SELSP) in English and Spanish. After measuring its parsing accuracy on test sets of these languages, we then apply our SESLP in the domain of SA for the polarity prediction task through a sentiment analysis engine which relies on syntax-based rules to calculate polarity in English and Spanish. (2) We compare the performance of our approach in terms of runtime and accuracy in both languages using the same syntactic rules on top of the well-known dependency parser from Stanza, as well as to VADER, a heuristic approach to SA that does not use syntax trees and is widely used in industry, and to a state-of-the-art Transformer-based (RoBERTa) model. (3) We study the influence of sentiment dictionaries and their interaction with syntax-based systems by using several sentiment dictionaries, and combinations thereof, on top of the same set of rules and the two mentioned parsers.
Related Work
@@
ML/DL methods and hybrid methods
Supervised approaches using machine learning, including deep learning architectures like LSTM or CNN and pre-trained models like BERT and RoBERTa, are a common practice in SA (see Cui et al., (2023)). These approaches use the standard supervised learning methodology where a training dataset is used during the training process to learn correlations between the specific input text and the sentiment polarity, and the resulting trained system is then evaluated on a separate test dataset and/or used in production. The performance of ML/DL-based approaches is known to be relatively high if the learning and prediction tasks are performed on a similar dataset or corpus, as they adapt especially well to the particularities of the training dataset. This means that they often dominate the rankings in shared tasks, such as that of Rest-Mex 2023 (Álvarez-Carmona et al.,, 2023), where training and test datasets are obtained from the same source and exhibit similar characteristics in terms of topic, length, style, etc. However, their performance can degrade on texts that differ from the training dataset, even within the same domain (see e.g. Table 5 of Vilares et al., (2017)).
@@
Some approaches, like (Yang et al.,, 2020), use sentiment dictionaries to provide or modify the input information fed to supervised machine learning approaches. Thus, they are often considered hybrid approaches (Cui et al.,, 2023) between ML/DL methods and the lexicon-based methods described in the next section. While exploiting sentiment dictionaries is useful to improve accuracy, their machine-learning-based core means that they still share the drawbacks described above.
Lexicon-based methods
@@
The lexicon-based methodology makes use of a sentiment dictionary that contains sentiment words with corresponding polarity scores to find out the overall opinion or polarity of a sentence or text (Taboada et al.,, 2011; Hutto and Gilbert,, 2014; Vilares et al.,, 2017; Kellert et al.,, 2023). These approaches fall into various branches depending on how they deal with syntactic rules that can shift or intensify the polarity. Non-compositional and non-recursive rule-based approaches such as SO-CAL (Taboada et al.,, 2011), SentiStrength (Thelwall et al.,, 2012) or VADER (Hutto and Gilbert,, 2014) use heuristic rules to account for polarity changes due to negation, modification, or other syntactic processes. Other unsupervised approaches use a more linguistically-informed architecture to account for SA by exploiting general syntactic rules that influence sentiment polarity, requiring a parse tree (Tan et al.,, 2012; Vilares et al.,, 2017; Kellert et al.,, 2023).
@@
Lexicon-based methods are considered to be simple to use and cost-effective as there is no need to train on huge amounts of data, and they work well across domains. For these reasons, in spite of less attention in the NLP literature in recent years due to the general trend towards machine learning in the field, they still tend to be the most commonly used in industry and production domains such as biomedicine (see e.g. Table 5 in (Denecke and Reichenpfader,, 2023)), fintech  (Bredice et al.,, 2025) or economics (Algaba et al.,, 2020; Barbaglia et al.,, 2025).
@@
To maximize accuracy, lexicon-based methods need to be able to reliably determine the scope of various relevant syntactic phenomena such as negation or intensification. As mentioned above, the most basic form of scope detection is to use simple local heuristics such as a fixed-length window, but this has limitations in terms of accuracy. For example, systems like SentiStrength or VADER can detect that “good” is negated in “not good”, but not in “not in any respect good”, due to the long-distance dependency. Instead, one can use machine learning classifiers to determine scope (see (Cruz et al.,, 2016; Jiménez-Zafra et al.,, 2021) and references therein), but this entails the disadvantages previously discussed for ML approaches, including the need for specific data and lack of explainability. Thus, the alternative on which we focus our proposal is to use syntactic rules that operate on a parse tree (Jia et al.,, 2009; Vilares et al.,, 2015, 2017; Jimenez Zafra et al.,, 2019; Kellert et al.,, 2023). The next sections address methods for syntactic parsing and how they can be used for polarity classification.
Syntactic parsing for sentiment analysis
@@
Syntactic parsing is the task of determining the hierarchical structure of a sentence in natural language. For this purpose, two main kinds of representations are used in NLP: constituent parsing, where the sentence is recursively divided into smaller units, and dependency parsing, which focuses on representing relations between words. While constituent parsing has seen some use in sentiment analysis (Socher et al.,, 2013), dependency parsing is the dominant alternative for the task, as it straightforwardly represents what words are modified by other words. In particular, it is the basis of all the methods based on syntactic rules cited in the previous section.
@@
In the last few years, Universal Dependencies (UD) (Zeman et al.,, 2023) has become the predominant framework for dependency parsing, especially in multilingual setups, as it provides harmonized annotation guidelines and freely-available corpora for over 100 diverse languages.
@@
Kellert et al., (2023) make use of UD-based syntactic parsing and exploit dependency relations expressed in UD between sentiment words and other words that can shift, strengthen, or weaken the polarity of sentiment words (see also Vilares et al., (2017) for application of syntactic parsing for SA using a predecessor framework to UD). The first stage of the approach in (Kellert et al.,, 2023) is to search for sentiment words taken from a sentiment dictionary SO-CAL (Brooke et al.,, 2009) in the input text, such as the sentiment word wonderful (in Figure 3). Then, the system traverses the dependency tree and applies a set of syntax-based rules to check whether negation words or other modifiers and relevant syntactic phenomena change the polarity of sentiment words (see the dependency tree in Figure 3). This syntactic approach applies recursive and compositional rules to negation and intensification, to identify the right scope of polarity shifting elements and to properly compute the sentence polarity (Vilares et al.,, 2017; Kellert et al.,, 2023). This means that the computing of the sentence polarity is done bottom-up: first, intensifiers and polarity-shifting syntactic phenomena in the structurally lowest levels of the tree are used to modify the polarity of their modified sentiment-bearing words and expressions. Then, structurally higher polarity shifting elements are computed, and polarities from converging branches of the dependency tree are aggregated, until the root of the sentence structure is reached and the final combined polarity for the whole sentence is obtained (Vilares et al.,, 2017; Kellert et al.,, 2023). For instance, in the sentence in Figure 3, first, the modification of the adjective “kind” by the adverb “very” is computed, and then the conjunction between the sentiment word “wonderful” and the output of “very kind” is computed. Since there are no further polarity-shifting elements, the polarity of this conjunction is output as sentence polarity.
Dependency parsing approaches
@@
To successfully use UD-based dependency parsing for sentiment analysis as described in the previous subsection, one ideally would want the parser to be accurate, so the obtained trees provide reliable information to compute polarity, and efficient, so that the system can operate at a large scale with a reasonable resource consumption. However, choosing a parsing approach involves a tradeoff between speed and accuracy, as the most accurate systems tend to be slower (Anderson and Gómez-Rodríguez,, 2021). In this respect, it is worth noting that Gómez-Rodríguez et al., (2019) showed that the polarity classification task has modest accuracy requirements, i.e., there is a “good enough” parsing accuracy threshold past which increasing accuracy doesn’t bring further improvements to the end task. On the other hand, speed is often a practical bottleneck, especially if a large number of messages need to be processed.
@@
However, as already noted in § Introduction, conventional syntactic parsers based on ad-hoc algorithms, like the one included in the Stanza library (Qi et al.,, 2020) are often slow and computationally costly (Gómez-Rodríguez and Vilares,, 2018; Roca et al.,, 2023; Strzyz et al., 2019b, ). However, in recent years, an alternative approach has been developed that provides considerably faster parsing, based on viewing dependency parsing as a sequence labeling task. Sequence labeling can be performed in linear time with respect to the length of the input, and is highly parallelizable, so it provides high practical speeds (Strzyz et al., 2019b, ; Anderson and Gómez-Rodríguez,, 2021).
@@
Previous work on syntactic parsing as sequence labeling has been applied to several NLP tasks such as integration of constituency and dependency parsing (Strzyz et al., 2019a, ), semantic role labeling (Wang et al.,, 2019) or LLM probing (Vilares et al.,, 2020), where its speed and simplicity can be useful. CoDeLin (COnstituent and DEpendency LINearization system, Roca et al., (2023)) is a recent comprehensive library for parsing as sequence labeling that supports both constituent and dependency parsing. The basic idea behind dependency parsing in CoDeLIn is to assign to each word a tag with information about a part of the dependency tree involving that word, in such a way that a full tree can be decoded from the sequence of labels for a given sentence. For this purpose, there are several types of encoding functions, each of which corresponds to a different way of representing the dependencies between words in a sentence in the form of tags associated to each of the words. We now summarize the framework and some of the encodings of interest. More information, including further encoding functions, can be found in (Strzyz et al., 2019b, ; Roca et al.,, 2023; Gómez-Rodríguez et al.,, 2023).
@@
Following (Roca et al.,, 2023), let $T_{|w|}$ be the set of dependency trees for sentences of length $|w|$. Each element of $T_{|w|}$ is a dependency tree $t_{w}=(V,A)$ for a sentence $w=[w_{1},...,w_{|w|}]$ where each $i\in V$ denotes the position of the word $w_{i}$, and $A$ is the set of dependency relations between words, where each $a\in A$ is of the form $a=(h,r,d)$ such that $h\in V$ is the index of the head of the relation, $d\in V$ corresponds to the index of the dependent, and $r\in R$ is a label describing a syntactic function, i.e. the type of syntactic relation between head and dependent (such as subject, modifier, indirect object, etc.). We can cast the dependency parsing task as a sequence labeling task by defining a set of labels $L$ that will let us encode each tree $t_{w}\in T_{|w|}$ as a sequence of $|w|$ labels $l_{w}\in L^{|w|}$, and an encoding function $F_{|w|}:T_{|w|}\rightarrow L^{|w|}$ that allows us to translate the trees into those labels. $F_{|w|}$ must be injective, so that each tree is encoded as a unique label sequence and correct label sequences can be decoded back into trees. While injectivity is not enough to guarantee that incorrect label sequences that might be output by a machine learning sequence-labeling system can also be decoded (that would require surjectivity, which known encoding functions do not have), such sequences are handled in practice with straightforward heuristics, like ignoring labels that are contradictory or violate tree constraints.
@@
Figure 1 illustrates how a sentence’s parse tree is represented in terms of different encodings in CoDeLIn (Strzyz et al., 2019b, ; Roca et al.,, 2023). Once one has chosen a tree encoding, training a parser is straightforward: a sequence-labeling model is trained to label sentences with the syntactic labels in the selected encoding, and at test time, the generated labels are decoded back to trees by using the displacements (and PoS tags, in the PoS-based encoding) to find the head of each node and build the corresponding dependency. If the model outputs a wrong label that causes a contradiction or constraint violation (e.g. a head offset pointing to a non-existing word position like position 6 in a sentence of length 4, or causing a cycle), that label is simply ignored and no dependency is created.
Figure 1: A short version for demonstration of Sequence Labeling from (Strzyz et al., 2019b, ).
@@
We are using different datasets for the different tasks in our pipeline. For dependency parsing, we use the datasets UD EWT English [^2] https://github.com/UniversalDependencies/UD_English-EWT/tree/master (Silveira et al.,, 2014) and UD Spanish AnCora [^3] https://github.com/UniversalDependencies/UD_Spanish-AnCora/tree/master (Taulé et al.,, 2008) with standard splits.